I have another word problem that I would like you to check.
a square piece of cardboard is formed into a box by cutting out 3 inch squares from each of the corners and folding up the sides.
if the volume of the box needs to be 867, what size square piece of cardboard is needed? I came up with a piece that is 20 x 20
867= (17)(17)(3) 20 x 20 - 3 inch corners =
17 x 17
would 20 x 20 be the correct answer?

Answer by checkley71(8403)

THE SIZE OF THE BOX ONCE THE CORNERS ARE CUT AND FOLDED IS NOW X~2*3=867 OR
X~2=867/3 OR X~2=289 OR X=17 THEREFORE THE ORIGINAL DIMENTIONS OF THE BOX HAD
TO BE 17+2*3 OR 17+6=23 INCHES SQUARE